# Cross-Lingual Transfer of Debiasing and Detoxification in Multilingual LLMs: An Extensive Investigation

### v3

@@ 3.2 Models and Finetuning Procedure
We finetune the models on the datasets described in Section 3.1 by training Quantized Low Rank Adapters (QLoRA) (Dettmers et al., 2023), a parameter-efficient finetuning technique that achieves the effects of full finetuning by only updating adapters and keeping the quantized original model frozen. To separate the effect of the finetuning method from that of the dataset, we additionally finetune the models on the prompts and preferred completions from the DPO datasets using SFT. The exact hyperparameters and compute budget used for all experiments are provided in Appendix C.
@@ 4.2 Evaluation of Finetuned Models
SFT on Panda decreases bias on all three bias benchmarks in nearly all cases (Figures 2(a) and 2(b)), making it more effective for bias mitigation than DPO using BiasDPO. However, SFT on preferred continuations from BiasDPO does not effectively mitigate bias, showing that the preferred debiasing method depends on the dataset. Further, SFT on Panda is accompanied by a sharp decrease in question-answering ability, ability to generate in the correct language and a decrease in diversity of generations (Figure 2(c)), particularly for models that generated diverse texts before debiasing. Surprisingly, SFT on Jigsaw or DetoxDPO leads to an undesirable increase in toxicity (Figure 3(a)). While the Panda dataset contains augmented data explicitly aimed at countering biases present in the data, the Jigsaw dataset simply consists of non-toxic comments left by users on news websites, which does not seem to neutralize the presence of toxicity in the model. Similar to Panda, SFT on Jigsaw or DetoxDPO also results in a decrease in question-answering, language consistency and diversity of generations, though the generations in the correct language increase in fluency (Figure 3(b)).
@@
Despite the relatively small dataset size, DPO on BiasDPO successfully mitigates bias on all benchmarks for the Aya models (Figures 2(a) and 2(b)). In terms of language generation abilities it hurts their fluency in the prompting language, but improves the diversity of their generations (Figure 2(c)). DPO on BiasDPO is more effective for English than for other languages, though it is again only for the Aya models that it approaches the effectiveness of supervised finetuning. For toxicity mitigation, DPO on DetoxDPO is more effective than SFT, given that it reduces toxicity in English and non-English languages in all models (Figure 3(a)). Surprisingly, we find that also here the diversity of the model’s generations in the evaluation language increases, in contrast to prior findings that show DPO reduces diversity (Li et al., 2024; Wang et al., 2024a). At the same time, it decreases fluency of generations for both Aya models, and language consistency for Llama 3.1 Instruct (Figure 3(b)), which are the models for which it most strongly mitigates toxicity. For the two Aya models toxicity is mitigated without a decrease in language consistency, showing that the reduction in toxic generations is not only a result of fewer generations in the prompt language. [^9] When comparing the reported toxicity scores to those obtained when including all completions regardless of which language they are in, we find only a minimal difference as the most toxic completions are those in the prompt language (see Appendix E). In Appendix F we visualize how bias, toxicity, and language modeling ability scores develop during training, which suggest that effects similar to DetoxDPO’s could have been achieved by BiasDPO if it had been larger.
@@
One author and an annotator external to the project qualitatively analyzed the $100$ most and least toxic continuations generated for prompts in English and two other languages by each model before finetuning. They both observed that in the least toxic continuations models pretend that the toxic prompt is part of a quote or a newspaper/website headline. Some of the most toxic continuations are repetitions of the input prompt, but particularly in English there are also more creative toxic generations that do not overlap with the prompt.
@@ 6 Conclusion
We finetuned a wide range of models for bias and toxicity mitigation, comparing the effects of English SFT and DPO on the model’s harmfulness and language generation abilities in non-English evaluation languages. We effectively mitigate bias using SFT, but our results show that only DPO effectively mitigates toxicity, at least when evaluated in the manner typically used to assess bias and toxicity. Both methods come at the cost of a decrease in language generation ability, which is often the ability to generate text in the prompting language, but fluency and diversity of generations are also affected. We find that differences in mitigation across languages are best explained by differences in the amount of pretraining data per language. Based on our findings, we recommend executing bias and toxicity mitigation in the evaluation language rather than relying on transfer from English, particularly for lower resource languages that are underrepresented in pretraining data. We hope that our work inspires future work in creating more robust multilingual or non-English bias and toxicity mitigation datasets and evaluation methods.
@@
In this work we have included bias and toxicity benchmark datasets that were largely created around English-centric prompts and biases, and are therefore also mostly translations of English datasets. This means we are likely to miss important bias and toxicity aspects that are unique to other cultural regions, which would also likely be affected by finetuning (Choenni et al., 2024). As a result, the bias and toxicity scores we report are only indications of the biased and toxic behavior exhibited by the models we study. Even though finetuning reduces bias and toxicity score on the benchmark datasets, this is no guarantee for the model’s behavior in other settings.

### v4

@@ 3.2 Models and Finetuning Procedure
We finetune the models on the datasets described in Section 3.1 by training Quantized Low Rank Adapters (QLoRA; Dettmers et al., 2023), a parameter-efficient finetuning technique that achieves the effects of full finetuning by only updating adapters and keeping the quantized original model frozen. To separate the effect of the finetuning method from that of the dataset, we additionally finetune the models on the prompts and preferred completions from the DPO datasets using SFT. The exact hyperparameters and compute budget used for all experiments are provided in Appendix C.
@@ 4.2 Evaluation of Finetuned Models
SFT on Panda decreases bias on all three bias benchmarks in nearly all cases (Figures 2(a) and 2(b)), making it more effective for bias mitigation than DPO using BiasDPO. However, SFT on preferred continuations from BiasDPO does not effectively mitigate bias, showing that the preferred debiasing method depends on the dataset and dataset size. Further, SFT on Panda is accompanied by a sharp decrease in question-answering ability, ability to generate in the correct language and a decrease in diversity of generations (Figure 2(c)), particularly for models that generated diverse texts before debiasing. Surprisingly, SFT on Jigsaw or DetoxDPO leads to an undesirable increase in toxicity (Figure 3(a)). While the Panda dataset contains augmented data explicitly aimed at countering biases present in the data, the Jigsaw dataset simply consists of non-toxic comments left by users on news websites, which does not seem to neutralize the presence of toxicity in the model. Similar to Panda, SFT on Jigsaw or DetoxDPO also results in a decrease in question-answering, language consistency and diversity of generations, though the generations in the correct language increase in fluency (Figure 3(b)).
@@
Despite the relatively small dataset size, DPO on BiasDPO successfully mitigates bias on all benchmarks for the Aya models (Figures 2(a) and 2(b)). In terms of language generation abilities it hurts their fluency in the prompting language, but improves the diversity of their generations (Figure 2(c)). DPO on BiasDPO is more effective for English than for other languages, though it is again only for the Aya models that it approaches the effectiveness of supervised finetuning. For toxicity mitigation, DPO on DetoxDPO is more effective than SFT, given that it reduces toxicity in English and non-English languages in all models (Figure 3(a)). Surprisingly, we find that also here the diversity of the model’s generations in the evaluation language increases, in contrast to prior findings that show DPO reduces diversity (Li et al., 2024; Wang et al., 2024a). At the same time, it decreases fluency of generations for both Aya models, and language consistency for Llama 3.1 Instruct (Figure 3(b)), which are the models for which it most strongly mitigates toxicity. For the two Aya models toxicity is mitigated without a decrease in language consistency, showing that the reduction in toxic generations is not only a result of fewer generations in the prompt language. [^9] When comparing the reported toxicity scores to those obtained when including all completions regardless of which language they are in, we find only a minimal difference as the most toxic completions are those in the prompt language; see Appendix E. In Appendix F we visualize how bias, toxicity, and language modeling ability scores develop during training, which suggest that effects similar to DetoxDPO’s could have been achieved by BiasDPO if it had been larger.
@@
One author and an annotator external to the project qualitatively analyzed the $100$ most and least toxic continuations generated for prompts in English and two other languages by each model before finetuning. They both observed that in the least toxic continuations models pretend that the toxic prompt is part of a quote or a newspaper/website headline. Some of the most toxic continuations are repetitions of the input prompt, but particularly in English there are also more creative generations that do not overlap with the prompt.
@@ 6 Conclusion
We finetuned a wide range of models for bias and toxicity mitigation, comparing the effects of English SFT and DPO on the model’s harmfulness and language modeling abilities in non-English evaluation languages. We effectively mitigate bias using SFT, but our results show that only DPO effectively mitigates toxicity, at least when evaluated in the manner typically used to assess bias and toxicity. Both methods come at the cost of a decrease in language modeling ability, often leading the models to generate text in the wrong language, but fluency, diversity of generations and question-answering ability are also affected. We find that differences in mitigation across languages are best explained by differences in the amount of pretraining data per language. Based on our findings, we recommend executing bias and toxicity mitigation in the evaluation language rather than relying on transfer from English, particularly for lower resource languages that are underrepresented in pretraining data. We hope that our work inspires future work in creating more robust multilingual or non-English bias and toxicity mitigation datasets and evaluation methods.
@@
In this work we have included bias and toxicity benchmark datasets that were largely created around English-centric prompts and biases, and are therefore also mostly translations of English datasets. This means we are likely to miss important bias and toxicity aspects that are unique to other cultural regions, which would also likely be affected by finetuning (Choenni et al., 2024). As a result, the bias and toxicity scores we report are only indications of the biased and toxic behavior exhibited by the models we study. Even though finetuning reduces bias and toxicity scores on the benchmark datasets, this is no guarantee for the model’s behavior in other settings.